Anna is four times as old as Matt, the sum of their ages is 65 years. how old are they
Found 2 solutions by ikleyn, stanbon:
Answer by ikleyn(52800)

.

x for Matt.


x + 4x = 65.


5x = 65.


You complete from this point.


Answer by stanbon(75887)

Anna is four times as old as Matt, the sum of their ages is 65 years. how old are they?
----
Symbolize with two equations
a = 4m
a+m = 65
----
Substitute for "a" and solve for "m"
4m+m = 65
5m = 65
m = 13 yrs (Matt's age)
------
Solve for "a"
a = 4m = 4*13 = 52 yrs (Anna's age)
----------------
Cheers,
Stan H.